# PointSeg: A Training-Free Paradigm for 3D Scene Segmentation via Foundation Models

### v4

@@ 1 Introduction
3D scene segmentation plays a vital role in many applications, such as autonomous driving, augmented reality and room navigation. To tackle the challenges in 3D scene segmentation, most of the previous methods (Kundu et al., 2020; Jiang et al., 2020; Rozenberszki et al., 2022; Kolodiazhnyi et al., 2024; Liang et al., 2021; Schult et al., 2023; Sun et al., 2023; Vu et al., 2022) are supervised and heavily rely on precise 3D annotations, which means that they lack the zero-shot capability. Despite recent efforts (Takmaz et al., 2023; Huang et al., 2023; Yin et al., 2023; He et al., 2024) have attempted to explore the zero-shot 3D scene understanding, these approaches either require 3D mask pre-trained networks or domain-specific data training. Consequently, the ability of domain transfer to unfamiliar 3D scenes continues to pose significant challenges.
@@
Looking around in the 2D realm, vision foundation models (VFMs) (Radford et al., 2021; Jia et al., 2021; Oquab et al., 2023; He et al., 2022) have exploded in growth, attributed to the availability of large-scale datasets and computational resources. And they have demonstrated exceptional generalization capabilities in zero-shot scenarios, along with multifunctional interactivity when combined with human feedback. Most recently, Segment Anything Model (SAM) (Kirillov et al., 2023) has managed to attain remarkable performance in class-agnostic segmentation by training on the SA-1B dataset. Then it triggers a series of applications in various tasks and improvements in various aspects (Zhang et al., 2023b; Zou et al., 2024; Xiong et al., 2023; Zhao et al., 2023; Zhang et al., 2023a). Inspired by this, a natural idea is to also train a foundation model in 3D space. However, this has been hindered by the limited scale of 3D data and the high cost of 3D data collection and annotation (Goyal et al., 2021; Chang et al., 2015). Considering this, we ask: Is it possible to explore the use of VFMs to effectively tackle a broad spectrum of 3D perception tasks without training, e.g., 3D scene segmentation ?
@@
Following this paradigm, some works have made some early attempts. One line focuses on segmenting 2D frame accurately with different scene deconstruction strategies (Yang et al., 2023b; Yin et al., 2024; Guo et al., 2023a). Another line tries to learn high-quality 3D points to prompt the SAM by using the projection from 3D to 2D (Xu et al., 2023). Though effective, none of these methods essentially acknowledge the facts of 3D scene segmentation in three challenging aspects: (i) 3D prompts are naturally prior to the one in the 2D space, which should be carefully designed rather than a simple projection, (ii) the initial segmentation mask from multiple views might include rough edges and isolated background noises, (iii) local adjacent frames maintain the global consensus, which might be overlooked during the merging process.
@@
To address these challenges, we present PointSeg, a novel perception framework that effectively incorporates different foundation models for tackling the 3D scene segmentation task without training. The key idea behind PointSeg is to learn accurate 3D point-box prompts pairs to enforce the off-the-shelf foundation models and fully unleash their potential in 3D scene segmentation tasks with three effective components. First, we construct a two-branch prompts learning structure to acquire the 3D point prompts and 3D box prompts respectively. The 3D point prompts are derived from localization abilities of PointLLM (Xu et al., 2024) to provide more explicit prompts in the form of points and 3D box prompts come from the 3D detectors (Shen et al., 2024; Wu et al., 2023b). Considering that the naive prompts could result in fragmented false-positive masks caused by matching outliers, we propose the bidirectional matching strategy for the generation of accurate point-box prompt pairs. Furthermore, when incorporated with different 2D vision segmentation foundation models, such as SAM 2 (Ravi et al., 2024), our approach involves the iterative post-refinement to eliminate the coarse boundaries and isolated instances of background noise adaptively. Finally, with the primary aim of segmenting all points within the 3D scene, we employ the affinity-aware merging algorithm to capture pairwise similarity scores based on the 2D masks generated by the vision segmentation foundation models.
@@
Comprehensive experiments on ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), and KITTI-360 (Liao et al., 2022) demonstrate the superior generalization of the proposed PointSeg, surpassing previous specialist training-free model by 14.9-16.3 mAP and specialist training-based methods by 5.6-8 mAP across different datasets, all without training on domain-specific data. Remarkably, our zero-shot approach yields superior results in comparison to fully-supervised PointSAM (Zhou et al., 2024) trained on synthetic datasets, thereby emphasizing the effectiveness of PointSeg in the segmentation of intricate 3D scene. Moreover, we incorporate different foundation models, i.e., SAM 2 (Ravi et al., 2024), SAM (Kirillov et al., 2023), FastSAM (Zhao et al., 2023), MobileSAM (Zhang et al., 2023a), and EfficientSAM (Xiong et al., 2023), into our pipeline, and the performance gain shows that enhancements on 2D images can be seamlessly translated to improve 3D results. We summarize the contributions of our paper as follows:
@@ 2 Related Work
Considering the point clouds in 3D space, 3D semantic segmentation task aims to predict a specific category towards the given point (Graham et al., 2018; Hu et al., 2020; Kundu et al., 2020; Li et al., 2018; Qi et al., 2017; Rozenberszki et al., 2022; Wang et al., 2019; Xu et al., 2018; Wang et al., 2015; Zhang et al., 2023c). 3D instance segmentation task broadens this concept by pinpointing separate entities within the same semantic class and bestowing unique masks upon each object instance (Choy et al., 2019; Fan et al., 2021; Han et al., 2020; Hou et al., 2019; Engelmann et al., 2020; Hui et al., 2022; Jiang et al., 2020; Kolodiazhnyi et al., 2024; Liang et al., 2021; Schult et al., 2023; Sun et al., 2023; Vu et al., 2022; Lahoud et al., 2019; Yang et al., 2019). Among them, Mask3D (Schult et al., 2023) designs a transformer-based network to build the 3D segmentation network and achieves the state-of-the-art performance. Although Mask3D has made significant progress, like previous supervised learning methods, it still necessitates a substantial volume of 3D annotated data for network training. This limitation impedes the method’s generalization to open-world scenarios that include new objects from unseen categories. Furthermore, the collection of annotated 3D data is not only costly but sometimes unfeasible due to privacy concerns. Our framework, however, aspires to directly leverage the intrinsic zero-shot potential of SAM for 3D scene segmentation, thereby negating the necessity for further model training.
Open-set 3D Segmentation.
@@
Inspired by the success of 2D open-vocabulary segmentation methods (Ghiasi et al., 2022; Liang et al., 2023), a series of works (Ding et al., 2023; Huang et al., 2023; Takmaz et al., 2023; Peng et al., 2023; He et al., 2024; Jiang et al., 2022) have dived to explore the potential of 3D open-vocabulary scene understanding. OpenMask3D (Takmaz et al., 2023) predicts 3D instance masks with the per-mask feature representations, which can be used for querying instances based on open-vocabulary concepts. OpenIns3D (Huang et al., 2023) employs a Mask-Snap-Lookup scheme to learn class-agnostic mask proposals and generate synthetic scene-level images at multiple scales. On the other hand, the interpolation capabilities of NeRFs (Mildenhall et al., 2021) are applied to integrate language with the CLIP feature by LERF (Kerr et al., 2023) and DFF (Kobayashi et al., 2022). And OR-NeRF (Yin et al., 2023) empowers users to segment an object by clicking and subsequently eliminate it from the scene. Although they have achieved encouraging instance segmentation results on indoor scenes with objects similar to the training data, these methods demonstrate a failure in complex scenes with fine-grained objects. In this study, we eliminate the reliance on a pre-trained 3D mask proposal network and instead focus directly on how to leverage the segmentation results of SAM to generate fine-grained 3D masks for 3D scenes.
Segment Anything Model in 3D.
@@
The emergence of the Segment Anything Model (Kirillov et al., 2023; Ravi et al., 2024) have triggered a revolution in the field of 2D segmentation. Having been trained on the extraordinary SA-1B dataset, SAM has garnered a vast amount of knowledge, equipping it to effectively segment unfamiliar images without additional training. Followed by SAM, several works (Zhang et al., 2023b; Zou et al., 2024; Xiong et al., 2023; Zhao et al., 2023; Zhang et al., 2023a; Liu et al., 2023) have attempted to accelerate or customize the original SAM from different aspects. Recognizing the exceptional capabilities of SAM, various recent research initiatives are working diligently to incorporate SAM into 3D scene segmentation task. Several works (Yang et al., 2023b; Yin et al., 2024; Guo et al., 2023a) attempt to segment each frame individually or constructs a graph based on the superpoints to lift the segmentation results to 3D space. However, these methods designate pixel prompts that are specific to each frame but do not synchronize across frames. This causes inconsistencies in segmentation across frames and produces substandard 3D segmentation results. Different from these 2D-to-3D lifting methods, SAMPro3D (Xu et al., 2023) attempts to locate 3D points in scenes as 3D prompts to align their projected pixel prompts across frames. Albeit effective, simply connecting 3D points to 2D space through projection is still too rough for complex scenes. In this paper, we propose a two-branch prompts learning structure towards accurate 3D prompts generation.
@@ 3 Method
As illustrated in Figure 2, we build a training-free 3D scene segmentation framework based on the off-the-shelf foundation models. Our PointSeg consists of three parts: 1) Bidirectional Matching based Prompts Generation (BMP) (Section 3.1). Given the reconstructed scene point cloud together with a set of posed RGB-D images , PointSeg first employs a two-branch prompts learning structure to acquire the 3D mask and box prompts, respectively. And the final point-box prompts pairs are obtained by the further bidirectional matching. These prompts serve as inputs to 2D vision foundation models, such as SAM 2 (Ravi et al., 2024), after aligning with the pixels in 2D images. 2) Iterative 2D Mask Generation (Section 3.2). Then, we perform Iterative Post-refinement (IPR) to enable the generation of mask proposals adaptively. 3) Affinity-aware 3D Mask Refinement (Section 3.3). We calculate the affinity scores between the points generated in the point-box pairs and the mask proposals, followed by the Affinity-aware Merging (AM) (see Algorithm 2) to obtain the final 3D segmentation masks.
@@ 3.1 Bidirectional Matching based Prompts Generation
Two-branch Prompts Generation.
Inspired by the strong ability of semantic understanding of some works (Xu et al., 2024; Zhang et al., 2022; Zhu et al., 2023) in 3D scene, we first employ PointLLM (Xu et al., 2024) for localization and rough segmentation in the upper branch in BMP of Figure 2. Given the point cloud , we also apply realistic projection to generate the different views, using the zero-initialized 3D grid , where / denote the spatial resolutions and represents the depth dimension vertical to the view plane. For each view, the normalized 3D coordinates of the input point cloud in a voxel in the grid can be denoted as
where is the scale factor to adjust the projected shape size. Following PointCLIPv2, we further apply the quantize, densify, smooth, and squeeze operations to obtain the projected depth maps . For the textual input, we utilize the large-scale language models (Brown et al., 2020) to obtain a series of 3D-specific descriptions. After feeding the depth maps and texts into their respective encoders, we can obtain the dense visual features where and the text feature . Then, we segment different parts of the shape on multi-view depth maps by dense alignment for each view and average the back-projected logits of different views into the 3D space, formulated as:
@@
Bidirectional Matching.
Given the coarse segmentation mask and the bounding box, we can already conduct the alignment to 2D images. However, the naive prompts often result in inaccurate and fragmented outcomes, riddled with numerous outliers. Therefore, we design the bidirectional matching strategy to put constraints on the point and box for high quality promptable segmentation.
@@
where denotes the cosine similarity between -th mask feature and -th box feature. Ideally, the matched regions should have the highest similarity. Then as illustrated in Figure 3, we propose to eliminate the matching outliers based on the similarity scores in three steps:
@@ 3.2 Iterative 2D Mask Generation
The vanilla segmentation model, such as SAM 2 (Ravi et al., 2024), accepts various inputs such as pixel coordinates, bounding boxes or masks and predict the segmentation area associated with each prompt. Hence, we feed the projected 2D point-box pairs calculated before into the foundation segmenters.
Iterative Post-refinement.
Through the above operation, we can obtain 2D segmentation mask on all frames from the decoder, which however, might include rough edges and isolated background noises. For further refinement, we iteratively feed the mask back into the decoder for the adaptive post-processing. As illustrated in Algorithm 1, we first obtain the 2D mask by feeding the 2D point-box pairs into the SAM-based decoder. Then we prompt the decoder additionally with this mask along with these projected prompt pairs to obtain the next mask. And the initial value to record change is set to infinity. In each subsequent iteration, we calculate the change ratio between the two adjacent masks and compare it with our predefined threshold , which is set to 5 by default. We repeat this iterative process until the change value falls below the threshold adaptively. It is worth noting that we have also tried the method of fixed iteration steps, but this adaptive iteration method have proved to be more effective which is shown in Figure 4 and Table 4.3.
@@ 3.3 Affinity-aware 3D Mask Refinement
Affinity-aware Merging.
Based on the -th projected point-box pair on the -th image and their 2D image segmentation mask, we compute the normalized distribution of the mask labels, denoted as . The affinity score between two projected points in the -th image can be computed by the cosine similarity between the two distributions, which can be represented as:
@@
Method Type ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 With Training SAM-graph (Guo et al., 2023a) specialist model 15.1 33.3 59.1 12.9 25.3 43.6 14.7 28.0 43.2 Mask3D (Schult et al., 2023) specialist model 26.9 44.4 57.5 8.8 15.0 22.3 0.1 0.4 4.2 OpenDAS (Yilmaz et al., 2024) specialist model 28.3 49.6 66.2 19.2 35.5 52.6 20.1 32.4 52.2 PointSAM (Zhou et al., 2024) specialist model 32.9 56.4 72.5 25.8 38.0 59.3 25.1 38.4 56.2 Training-free SAM3D (Yang et al., 2023b) specialist model 13.7 29.7 54.5 8.3 17.5 33.7 6.3 16.0 35.6 SAI3D (Yin et al., 2024) specialist model 18.8 42.5 62.3 17.1 31.1 49.5 16.5 30.2 48.6 SAMPro3D (Xu et al., 2023) specialist model 22.2 45.6 65.7 18.9 33.7 51.6 17.3 31.1 49.6 PointSeg (Ours) generalist model 38.5 63.6 82.1 33.8 49.1 67.2 32.3 47.2 66.5
Method ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 + SAM 2 (Ravi et al., 2024) 38.5 63.6 82.1 33.8 49.1 67.2 32.3 47.2 66.5 + SAM (Kirillov et al., 2023) 36.3 60.2 79.3 31.2 46.5 64.8 29.9 44.5 63.3 + MobileSAM (Zhang et al., 2023a) 26.2 49.8 68.3 19.6 36.4 55.2 20.6 34.6 53.3 + FastSAM (Zhao et al., 2023) 26.9 50.8 69.1 20.5 37.7 56.5 21.2 35.8 54.5 + EfficientSAM (Xiong et al., 2023) 33.5 57.2 75.8 27.8 43.6 62.4 27.5 41.7 60.6

### v5

@@ 1 Introduction
3D scene segmentation plays a vital role in many applications, such as autonomous driving, augmented reality and room navigation. To tackle the challenges in 3D scene segmentation, most of the previous methods [27, 21, 41, 26, 32, 42, 44, 46] are supervised and heavily rely on precise 3D annotations, which means that they lack the zero-shot capability. Despite recent efforts [45, 18, 63, 15] have attempted to explore the zero-shot 3D scene understanding, these approaches either require 3D mask pre-trained networks or domain-specific data training. Consequently, the ability of domain transfer to unfamiliar 3D scenes continues to pose significant challenges.
@@
Looking around in the 2D realm, vision foundation models (VFMs) [39, 20, 36, 14] have exploded in growth, attributed to the availability of large-scale datasets and computational resources. And they have demonstrated exceptional generalization capabilities in zero-shot scenarios, along with multifunctional interactivity when combined with human feedback. Most recently, Segment Anything Model (SAM) [24] has managed to attain remarkable performance in class-agnostic segmentation by training on the SA-1B dataset. Then it triggers a series of applications in various tasks and improvements in various aspects [67, 72, 52, 69, 65]. Inspired by this, a natural idea is to also train a foundation model in 3D space. However, this has been hindered by the limited scale of 3D data and the high cost of 3D data collection and annotation [9, 2]. Considering this, we ask: Is it possible to explore the use of VFMs to effectively tackle a broad spectrum of 3D perception tasks without training, e.g., 3D scene segmentation?
@@
Following this paradigm, some works have made some early attempts. One line focuses on segmenting 2D frame accurately with different scene deconstruction strategies [59, 64, 11]. Another line tries to learn high-quality 3D points to prompt the SAM by using the projection from 3D to 2D [53]. Though effective, none of these methods essentially acknowledge the facts of 3D scene segmentation in three challenging aspects: (i) 3D prompts are naturally prior to the one in the 2D space, which should be carefully designed rather than a simple projection, (ii) the initial segmentation mask from multiple views might include rough edges and isolated background noises, (iii) local adjacent frames maintain the global consensus, which might be overlooked during the merging process.
@@
To address these challenges, we present PointSeg, a novel perception framework that effectively incorporates different foundation models for tackling the 3D scene segmentation task without training. The key idea behind PointSeg is to learn accurate 3D point-box prompts pairs to enforce the off-the-shelf foundation models and fully unleash their potential in 3D scene segmentation tasks with three effective components. First, we construct a two-branch prompts learning structure to acquire the 3D point prompts and 3D box prompts respectively. The 3D point prompts are derived from localization abilities of PointLLM [54] to provide more explicit prompts in the form of points and 3D box prompts come from the 3D detectors [43, 51]. Considering that the naive prompts could result in fragmented false-positive masks caused by matching outliers, we propose the bidirectional matching strategy for the generation of accurate point-box prompt pairs. Furthermore, when incorporated with different 2D vision segmentation foundation models, such as SAM 2 [40], our approach involves the iterative post-refinement to eliminate the coarse boundaries and isolated instances of background noise adaptively. Finally, with the primary aim of segmenting all points within the 3D scene, we employ the affinity-aware merging algorithm to capture pairwise similarity scores based on the 2D masks generated by the vision segmentation foundation models.
@@
Comprehensive experiments on ScanNet [4], ScanNet++ [61], and KITTI-360 [33] demonstrate the superior generalization of the proposed PointSeg, surpassing previous specialist training-free model by 14.9-16.3 mAP and specialist training-based methods by 5.6-8 mAP across different datasets, all without training on domain-specific data. Remarkably, our zero-shot approach yields superior results in comparison to fully-supervised PointSAM [70] trained on synthetic datasets, thereby emphasizing the effectiveness of PointSeg in the segmentation of intricate 3D scene. Moreover, we incorporate different foundation models, i.e., SAM 2 [40], SAM [24], FastSAM [69], MobileSAM [65], and EfficientSAM [52], into our pipeline, and the performance gain shows that enhancements on 2D images can be seamlessly translated to improve 3D results. We summarize the contributions of our paper as follows:
@@ 2 Related Work
Closed-set 3D Segmentation. Considering the point clouds in 3D space, 3D semantic segmentation task aims to predict a specific category towards the given point [10, 17, 27, 30, 38, 41, 49, 55, 48, 68]. 3D instance segmentation task broadens this concept by pinpointing separate entities within the same semantic class and bestowing unique masks upon each object instance [3, 7, 13, 16, 6, 19, 21, 26, 32, 42, 44, 46, 28, 58]. Among them, Mask3D [42] designs a transformer-based network to build the 3D segmentation network and achieves the state-of-the-art performance. Although Mask3D has made significant progress, like previous supervised learning methods, it still necessitates a substantial volume of 3D annotated data for network training. This limitation impedes the method’s generalization to open-world scenarios that include new objects from unseen categories. Furthermore, the collection of annotated 3D data is not only costly but sometimes unfeasible due to privacy concerns. Our framework, however, aspires to directly leverage the intrinsic zero-shot potential of SAM for 3D scene segmentation, thereby negating the necessity for further model training.
@@
Open-set 3D Segmentation. Inspired by the success of 2D open-vocabulary segmentation methods [8, 31], a series of works [5, 18, 45, 37, 15, 22] have dived to explore the potential of 3D open-vocabulary scene understanding. OpenMask3D [45] predicts 3D instance masks with the per-mask feature representations, which can be used for querying instances based on open-vocabulary concepts. OpenIns3D [18] employs a Mask-Snap-Lookup scheme to learn class-agnostic mask proposals and generate synthetic scene-level images at multiple scales. On the other hand, the interpolation capabilities of NeRFs [35] are applied to integrate language with the CLIP feature by LERF [23] and DFF [25]. And OR-NeRF [63] empowers users to segment an object by clicking and subsequently eliminate it from the scene. Although they have achieved encouraging instance segmentation results on indoor scenes with objects similar to the training data, these methods demonstrate a failure in complex scenes with fine-grained objects. In this study, we eliminate the reliance on a pre-trained 3D mask proposal network and instead focus directly on how to leverage the segmentation results of SAM to generate fine-grained 3D masks for 3D scenes.
@@
Segment Anything Model in 3D. The emergence of the Segment Anything Model [24, 40] have triggered a revolution in the field of 2D segmentation. Having been trained on the extraordinary SA-1B dataset, SAM has garnered a vast amount of knowledge, equipping it to effectively segment unfamiliar images without additional training. Followed by SAM, several works [67, 72, 52, 69, 65, 34] have attempted to accelerate or customize the original SAM from different aspects. Recognizing the exceptional capabilities of SAM, various recent research initiatives are working diligently to incorporate SAM into 3D scene segmentation task. Several works [59, 64, 11] attempt to segment each frame individually or constructs a graph based on the superpoints to lift the segmentation results to 3D space. However, these methods designate pixel prompts that are specific to each frame but do not synchronize across frames. This causes inconsistencies in segmentation across frames and produces substandard 3D segmentation results. Different from these 2D-to-3D lifting methods, SAMPro3D [53] attempts to locate 3D points in scenes as 3D prompts to align their projected pixel prompts across frames. Albeit effective, simply connecting 3D points to 2D space through projection is still too rough for complex scenes. In this paper, we propose a two-branch prompts learning structure towards accurate 3D prompts generation.
@@ 3 Method
As illustrated in Fig. 2, we build a training-free 3D scene segmentation framework based on the off-the-shelf foundation models. Our PointSeg consists of three parts: 1) Bidirectional Matching based Prompts Generation (BMP) (Section 3.1). Given the reconstructed scene point cloud together with a set of posed RGB-D images , PointSeg first employs a two-branch prompts learning structure to acquire the 3D mask and box prompts, respectively. And the final point-box prompts pairs are obtained by the further bidirectional matching. These prompts serve as inputs to 2D vision foundation models, such as SAM 2 [40], after aligning with the pixels in 2D images. 2) Iterative 2D Mask Generation (Section 3.2). Then, we perform Iterative Post-refinement (IPR) to enable the generation of mask proposals adaptively. 3) Affinity-aware 3D Mask Refinement (Section 3.3). We calculate the affinity scores between the points generated in the point-box pairs and the mask proposals, followed by the Affinity-aware Merging (AM) (see Algorithm 2) to obtain the final 3D segmentation masks.
@@ 3.1 Bidirectional Matching based Prompts Generation
Two-branch Prompts Generation. Inspired by the strong ability of semantic understanding of some works [54, 66, 71] in 3D scene, we first employ PointLLM [54] for localization and rough segmentation in the upper branch in BMP of Figure 2. Given the point cloud , we also apply realistic projection to generate the different views, using the zero-initialized 3D grid , where / denote the spatial resolutions and represents the depth dimension vertical to the view plane. For each view, the normalized 3D coordinates of the input point cloud in a voxel in the grid can be denoted as
where is the scale factor to adjust the projected shape size. Following PointCLIPv2, we further apply the quantize, densify, smooth, and squeeze operations to obtain the projected depth maps . For the textual input, we utilize the large-scale language models [1] to obtain a series of 3D-specific descriptions. After feeding the depth maps and texts into their respective encoders, we can obtain the dense visual features where and the text feature . Then, we segment different parts of the shape on multi-view depth maps by dense alignment for each view and average the back-projected logits of different views into the 3D space, formulated as:
@@
Bidirectional Matching. Given the coarse segmentation mask and the bounding box, we can already conduct the alignment to 2D images. However, the naive prompts often result in inaccurate and fragmented outcomes, riddled with numerous outliers. Therefore, we design the bidirectional matching strategy to put constraints on the point and box for high quality promptable segmentation.
@@
where denotes the cosine similarity between -th mask feature and -th box feature. Ideally, the matched regions should have the highest similarity. Then as illustrated in Fig. 3, we propose to eliminate the matching outliers based on the similarity scores in three steps:
@@
Input: the projected point coordinates and box in frame , the predefined threshold
Output: the refined mask
@@ 3.2 Iterative 2D Mask Generation
The vanilla segmentation model, such as SAM 2 [40], accepts various inputs such as pixel coordinates, bounding boxes or masks and predict the segmentation area associated with each prompt. Hence, we feed the projected 2D point-box pairs calculated before into the foundation segmenters.
Input: affinity matrix where indicates the affinity score between two points and , and is the number of points.
Output: mask label .
Iterative Post-refinement. Through the above operation, we can obtain 2D segmentation mask on all frames from the decoder, which however, might include rough edges and isolated background noises. For further refinement, we iteratively feed the mask back into the decoder for the adaptive post-processing. As illustrated in Algorithm 1, we first obtain the 2D mask by feeding the 2D point-box pairs into the SAM-based decoder. Then we prompt the decoder additionally with this mask along with these projected prompt pairs to obtain the next mask. And the initial value to record change is set to infinity. In each subsequent iteration, we calculate the change ratio between the two adjacent masks and compare it with our predefined threshold , which is set to 5 by default. We repeat this iterative process until the change value falls below the threshold adaptively. It is worth noting that we have also tried the method of fixed iteration steps, but this adaptive iteration method have proved to be more effective which is shown in Fig. 4 and Table 6.
@@ 3.3 Affinity-aware 3D Mask Refinement
Affinity-aware Merging. Based on the -th projected point-box pair on the -th image and their 2D image segmentation mask, we compute the normalized distribution of the mask labels, denoted as . The affinity score between two projected points in the -th image can be computed by the cosine similarity between the two distributions, which can be represented as: